3x2 +18x + 21= x(x + 5)
Answer by checkley77(12844)

3x^2 +18x + 21= x(x + 5)
3x^2+18x+21=x^2+5x
3x^2-x^2+18x-5x+21=0
2x^2+13x+21=0
(2x+7)(x+3)=0
2x+7=0
2x=-7
x=-3.5 ans.
x+3=0
x=-3 ans.